OpenStudy (anonymous):

tan of theta= - square root of 3 in radians

OpenStudy (anonymous):

is \(-\sqrt3\) in radians?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i was just being an retrice making stupid joke \(-\sqrt3\) is just a number as \(\theta\) is just a number too the "in radians" part is like saying "in Fahrenheit"

OpenStudy (anonymous):

in any case, since \[\tan(\theta)=\frac{\sin(\theta)}{\cos(\theta)}\] look for where \(\sin(\theta)=-\frac{\sqrt{3}}{2}\)and \(\cos(\theta)=\frac{1}{2}\) or where \(\sin(\theta)=\frac{\sqrt{3}}{2}\) and \(\cos(\theta)=-\frac{1}{2}\) there are two such points on the unit circle

OpenStudy (anonymous):

look at the unit circle on the last page of the attached cheat sheet you will see the coordinates \((-\frac{1}{2},\frac{\sqrt3}{2})\) there